### v3

# Improving Low-Resource Knowledge Tracing Tasks by Supervised Pre-training and Importance Mechanism Fine-tuning

Abstract
Knowledge tracing (KT) aims to estimate student’s knowledge mastery based on their historical interactions. Recently, the deep learning based KT (DLKT) approaches have achieved impressive performance in the KT task. These DLKT models heavily rely on the large number of available student interactions. However, due to various reasons such as budget constraints and privacy concerns, observed interactions are very limited in many real-world scenarios, a.k.a, low-resource KT datasets. Directly training a DLKT model on a low-resource KT dataset may lead to overfitting and it is difficult to choose the appropriate deep neural architecture. Therefore, in this paper, we propose a low-resource KT framework called LoReKT to address above challenges. Inspired by the prevalent “pre-training and fine-tuning” paradigm, we aim to learn transferable parameters and representations from rich-resource KT datasets during the pre-training stage and subsequently facilitate effective adaptation to low-resource KT datasets. Specifically, we simplify existing sophisticated DLKT model architectures with purely a stack of transformer decoders. We design an encoding mechanism to incorporate student interactions from multiple KT data sources and develop an importance mechanism to prioritize updating parameters with high importance while constraining less important ones during the fine-tuning stage. We evaluate LoReKT on six public KT datasets and experimental results demonstrate the superiority of our approach in terms of AUC and Accuracy. To encourage reproducible research, we make our data and code publicly available at https://anonymous.4open.science/r/LoReKT-C619.
keywords:
Educational data mining , Knowledge tracing , Pre-training and fine-tuning , Importance mechanism
[…]
1 Introduction
Knowledge tracing holds a pivotal role within the realm of Intelligent Tutoring Systems (ITS) [1, 2, 3, 4]. Its primary objective is to forecast students’ performance on questions by estimating their mastery of individual knowledge components (KCs [^1] A KC is a generality of everyday terms like concept, principle, or skill.) through an analysis of their past interactions. A KC is a description of a mental structure or process that a learner uses, alone or in combination with other KCs, to accomplish steps in a task or a problem. Take Figure 1 as an example. The student has successively responded to four questions ($Q_{1}$ to $Q_{4}$), achieving correct answers for $Q_{1}$ and $Q_{3}$, while $Q_{2}$ is answered incorrectly. This pattern suggests that the student may have a proficient understanding of the “Addition”, “Subtraction”, and “Multiplication” KCs, but lacks familiarity with the “Modulo” and “Division” KCs. Leveraging the current knowledge mastery, the KT task aims to predict the student’s performance on the upcoming sixth question, $Q_{4}$. After gaining insights into students’ knowledge mastery through KT, educators can promptly pinpoint weaknesses and provide targeted exercises for improvement. Additionally, this information can assist online learning platforms in providing a series of adaptive learning services such as learning resource recommendations, customizing student learning paths, and personalizing teaching plans [5, 6].
Recently, with the remarkable progress of deep learning techniques, many studies develop deep learning based KT (DLKT) models that are trained on massive students’ historical interactions to pursue high accuracy on students’ knowledge mastery estimations. Thus, many publicly available educational datasets have been released for training an effective DLKT model.
Figure 1: An illustration of the KT problem.
However, due to the users’ privacy protection of the educational applications and the different learning energy and enthusiasm of students, it is extremely difficult to collect large-scale high-quality student interaction sequences from real-world educational environments. Therefore, the educational datasets for KT model training frequently involve limited student learning records. For example, ASSISTments2009 is one of the classical KT datasets and the observed interaction records are only collected from 4,217 students. However, most of the state-of-the-art DLKT models are designed with stacks of neural networks such as recurrent neural networks, memory networks [7, 8, 9, 10]. Directly training a DLKT model on such a low-resource KT dataset is very easy to run into the problem of overfitting. Furthermore, it is unclear what type of model architectures are most suitable for low-resource KT datasets in previous KT works.
To enhance the learning capability from low-resource datasets of deep learning-based models, some studies perform “pre-training and fine-tuning” paradigm [11, 12, 13, 14]. This paradigm leverages rich-resource datasets to pre-train a model first and transfer the learned parameters to the low-resource dataset. Motivated by these promising studies, we propose a simple yet effective framework called LoReKT. LoReKT aims to improve the performance on the low-resource KT dataset by transferring the knowledge tracing capability from the model pre-trained on multiple rich-resource KT datasets. More specifically, in the pre-training stage, we build a foundational pre-trained KT model using a stack of transformer decoders based on multiple rich-resource KT datasets. To enhance the model’s capacity for integrating information from both questions and concepts, we introduce data type embeddings. Furthermore, to enable the model to learn the distinct and shared tracing patterns from multiple KT datasets, we introduce dedicated dataset embeddings for each KT dataset. In the fine-tuning stage, we propose an importance vector-based fine-tuning strategy to allow the model to focus on updating crucial parameters for the specific target low-resource dataset while constraining unimportant parameters to prevent the learning and memorization of noisy information.
As a result, the implementation of LoReKT has the following merits:
  • 1.
    The framework avoids direct training by initially pre-training the model on rich-resource KT datasets and subsequently fine-tuning it on specific low-resource KT datasets. This approach mitigates the risk of overfitting.
  • 2.
    The framework leverages a stack of transformer decoders as its backbone, which has demonstrated excellent performance in various “pre-training and fine-tuning” scenarios. Moreover, this backbone requires no additional architecture design efforts, simplifying the process and reducing the reliance on specific model architectures for particular datasets.
  • 3.
    To further mitigate the overfitting issue, the framework employs a fine-tuning strategy based on an importance mechanism, restricting the learning of less important parameters to prevent the memorization of noisy information.
  • 4.
    To ensure that our approach can be fairly comparable with other recently developed DLKT models, we follow a publicly available standardized KT task evaluation protocol [15]. We conduct comprehensive and rigorous experiments on three public rich-resource datasets and three low-resource datasets. The results show that after pre-training, the pre-trained KT model comes close to the performance of previous approaches on rich-resource datasets. By fine-tuning the pre-trained KT model on the low-resource datasets, it achieves superior prediction performance in terms of AUC and Accuracy compared to 17 baselines.
2 Related Work
2.1 Deep Learning based Knowledge Tracing
Deep Knowledge Tracing (DKT) has pioneered the application of deep learning in knowledge tracing tasks by employing a Long Short Term Memory (LSTM) layer to encapsulate students’ knowledge states and predict students’ response performances [16]. Since then, many methods tend to use deep learning techniques to solve KT problem [1, 2, 3, 4, 7, 17, 8, 18, 19, 20, 21, 22, 10, 23, 24]. For example, Yeung and Yeung [17] leveraged prediction-consistent regularization mechanism to mitigate issues related to input reconstruction failure and prediction inconsistency in the context of DKT [16]. Zhang et al. [7] integrated a meticulously designed static key matrix for storing the interconnections among different knowledge components (KCs). Simultaneously, it utilizes a dynamic value matrix to iteratively update the knowledge state of students. Motivated by the learning curve theory [25]Nagatani et al. [8] took student’s forgetting behavior into consideration to enhance DKT [16]. Lee and Yeung [18] used student knowledge state encoder and skill encoder to predict the student response performance via the dot product. Tiana et al. [21] performed multi-task learning based on the bidirectional encoder representations to construct mixed representations of questions. To mitigate the potential issue of limited generalization in DLKT, adversarial training techniques, such as adversarial perturbations, are introduced to the original student interaction sequence. Specifically, Guo et al. [10] improveed the generalization capability of the DLKT model by incorporating adversarial perturbations at the embedding level of the student interaction sequence. The carefully designed perturbations contribute to the model’s effective generalization across diverse student interactions. Moreover, certain studies have concentrated on exploring the interactions between student responses and questions, as well as the associations between questions and KCs. For example, Nakagawa et al. [26] constructed a question-concept knowledge graph and utilized graph neural network to aggregate the node features related to the corresponding concepts and subsequently updates the student’s knowledge states effectively. Additionally, Ma et al. [23] employed self-supervised learning paradigm to identify the latent relationship between questions and KCs, thereby enhancing input representations. Another research direction focused on the interdependence among student interactions, aiming to capture finer details embedded within them. For instance, Pandey and Karypis [19] utilized a self-attention mechanism to grasp the relationships between exercises and students’ responses. Ghosh et al. [20] presented AKT, which utilizes two self-attention modules to extract the inner relevance of questions and interactions respectively, and explicitly model students’ forgetting behaviors via a monotonic attention mechanism.
In this paper, unlike the aforementioned DLKT methods that are committed to developing a series of sophisticated architectures, our LoReKT is based on a stack of simple transformer decoders. This unified backbone aims to break down disciplinary barriers and learn consistent representations across multiple KT datasets.
2.2 Pre-training for Low-resource Setting
In real-world scenarios, encounters with low-resource settings are commonplace, and the paradigm of ”pre-training and fine-tuning” has consistently proven its efficacy in addressing challenges within such contexts [27, 28, 29, 30, 31, 32, 33]. For example, Yosinski et al. [27] explored the transferability of AlexNet and observed that the initial three layers of AlexNet encapsulate general features conducive to transferability. By introducing fine-tuning to the neural network, it successfully mitigated data variability and scarcity, consequently enhancing the overall network performance. Bansal et al. [29] introduced a straightforward methodology to enhance direct speech-to-text translation (ST) in scenarios where the source language is low-resource. The approach involves initial pre-training of the model on a high-resource automatic speech recognition (ASR) task, followed by a subsequent fine-tuning process to refine its parameters specifically for speech-to-text translation (ST). Zhang et al. [30] proposed an innovative approach, the adaptive data augmentation fine-tuning technique, designed to facilitate the efficient transfer of Named Entity Recognition (NER) knowledge from resource-rich domains to low-resource target domains. Liu et al. [31] utilized pre-training data in both the initial pre-training and subsequent fine-tuning stages, strategically enhancing the model’s performance across low-resource datasets. This dual-stage utilization of pre-training data contributes to a comprehensive and effective optimization, addressing the challenges posed by limited data availability in low-resource scenarios. Chi et al. [32] investigated the prospect of enhancing the performance of low-resource non-English languages by incorporating pre-trained language models that are primarily trained on English. This exploration seeks to leverage the knowledge embedded in English-dominant language models to boost the capabilities of models applied to non-English languages with limited resources. To generate high-quality definition for low-resource language, Zhang et al. [33] leveraged a multilingual pre-trained model as backbone and employed a prompt contrastive fine-tuning approach to enhance the model’s capabilities in this specific linguistic context.
In this paper, we adhere to the ”pre-training and fine-tuning” paradigm, opting for a strategic approach rather than direct training of a DLKT model on low-resource KT datasets. This choice is made to mitigate potential overfitting issues and enhance the overall performance of model in low-resource KT datasets scenarios.
3 Problem Statement
The objective of KT problem is to predict the probability of whether a student will answer arbitrary $q_{*}$ correctly based on the student’s historical interaction data. Specifically, suppose a student’s chronologically ordered collection of $T$ past interactions is denoted as $\mathbf{S}=\{\mathbf{s}_{j}\}_{j=1}^{T}$, each student interaction $\mathbf{s}_{j}$ is represented as an ordering 4-tuple, i.e., $\mathbf{s}_{j}=<q_{j},\{c|c\in\mathcal{N}_{q_{j}}\}_{j},r_{j},t_{j}>$, where $q_{j}$, $\{c\}_{j}$, $r_{j}$ and $t_{j}$ represent the specific question, the associated KC set, student response [^2] Response $r_{j}\in\{0,1\}$, 1 represents the student answered correctly, and 0 otherwise. and student’s response timestamp respectively. $\mathcal{N}_{q_{j}}$ is the set of KCs that are associated with the question $q_{j}$. We would like to estimate the probability $\hat{r}_{*}$ of the student’s future performance on arbitrary question $q_{*}$.
4 The Framework
In this section, we introduce the procedures in our proposed LoReKT framework in details: (1) obtaining a foundational pre-trained KT model through learning from rich-resource KT datasets during the pre-training stage (Section 4.1); (2) efficiently adapting the pre-trained KT model to the low-resource dataset using an importance vector in the fine-tuning stage (Section 4.2).
Figure 2: An illustration of the interaction encoding and forward procedure of our LoReKT framework.
4.1 Pre-training Stage
Our objective in the pre-training stage is to learn transferable parameters and representations from certain rich-resource datasets and build a pre-trained KT foundation model that is able to quickly adapt to low-resource KT datasets.
4.1.1 Interaction Encoding
Due to the fact that the question bank is typically much larger than KCs, previous research mainly used KCs for interaction encoding, treating questions with the same KCs as identical [16, 34, 35]. However, this approach led DLKT models to overlook the unique characteristics of same-KC questions, limiting interaction representation. To address this, we align with works that include both individual question features and KCs to encode interactions in a more granular manner [36, 20, 24].
Specifically, let $\mathbf{D}=\{D_{i}\}_{i=1}^{I}$ be the mixed students’ learning sequences, where $D_{i}=\{\mathbf{S}^{i}_{1},...,\mathbf{S}^{i}_{n}\}$. $\mathbf{S}^{i}_{j}$ is the $j$th learning sequence from rich-dataset $i$. $n$ is the number of learning sequences in $D_{i}$, and $I$ is the number of rich-resource KT datasets. Let $\mathcal{N}_{q_{t}}$ be the set of KCs associated with $q_{t}$. We represent question $q_{t}$ and its corresponding KCs as follows:
$\mathbf{q}_{t}=\mathbf{W}^{q}\cdot\mathbf{e}^{q}_{t}$
$\bar{\mathbf{c}}_{t}=\frac{1}{\left|C_{q_{t}}\right|}\sum_{j=1}^{M}\mathbf{c}_{j}*\mathbb{I}(c_{j}\in\mathcal{N}_{q_{t}})$ (1)
$\mathbf{c}_{j}=\mathbf{W}^{c}\cdot\mathbf{e}^{c}_{j}$
where $\mathbf{e}^{q}_{t}\in\mathbb{R}^{N\times 1}$ and $\mathbf{e}^{c}_{j}\in\mathbb{R}^{M\times 1}$ are the one-hot vectors that indicating the question $q_{t}$ and the related KC in $\mathcal{N}_{q_{t}}$. $\mathbf{c}_{j}\in\mathbb{R}^{d\times 1}$ is one of the latent representations of the related KC to question $q_{t}$. $\mathbf{q}_{t}\in\mathbb{R}^{d\times 1}$ and $\bar{\mathbf{c}}_{t}\in\mathbb{R}^{d\times 1}$ are the latent embedding of $q_{t}$ and its corresponding KCs, respectively. $\mathbf{W}^{q}\in\mathbb{R}^{d\times N}$ and $\mathbf{W}^{c}\in\mathbb{R}^{d\times M}$ are learnable linear transformation operations. $N$ and $M$ are the total number of distinct questions and KCs in our mixed dataset $\mathbf{D}$, respectively [^3] We reassign ID numbers for all questions and KCs in the dataset $\mathbf{D}$ based on the values of $N$ and $M$. In the fine-tuning stage, for specific low-resource KT dataset, we adjust the ID numbers for its questions and KCs, starting from $N$ and $M$. Additionally, we expand the size of $\mathbf{W}^{q}$ and $\mathbf{W}^{c}$ to obtain their corresponding question and KC representations, i.e., $\mathbf{q}_{t}$ and $\mathbf{c}_{j}$.. $\cdot$ is the standard matrix/vector multiplication. $C_{q_{t}}$ is the size of $\mathcal{N}_{q_{t}}$ and $\mathbb{I}(\cdot)$ is the indicator function.
Drawing inspiration from the powerful pre-trained model BERT [37], which leverages token type embeddings to improve the integration of various token information, we introduce data type embeddings to the KT problem. In this problem, there are two distinct data types: questions and concepts. To improve the pre-trained KT model’s ability to incorporate information from both, we introduce question and concept data type embeddings, which are directly integrated into all question and concept embeddings:
$\widetilde{\mathbf{q}}_{t}=\mathbf{q}_{t}\oplus\mathbf{t}_{q};\quad\widetilde{\mathbf{c}}_{t}=\bar{\mathbf{c}}_{t}\oplus\mathbf{t}_{c}$ (2)
where $\mathbf{t}_{q}\in\mathbb{R}^{d\times 1}$ and $\mathbf{t}_{c}\in\mathbb{R}^{d\times 1}$ are the question and concept data type embeddings, $\oplus$ is the element-wise addition operator. $\widetilde{\mathbf{q}}_{t}$ and $\widetilde{\mathbf{c}}_{t}$ are the question and KCs embeddings enriched with data type information.
Finally, we combine the embedding of question, its corresponding KCs, and response to encode the interaction $\mathbf{e}_{t}\in\mathbb{R}^{d\times 1}$, i.e.:
$\mathbf{x}_{t}=\widetilde{\mathbf{q}}_{t}\oplus\widetilde{\mathbf{c}}_{t};\quad\mathbf{r}_{t}=\mathbf{W}^{a}\cdot\mathbf{a}^{q}_{t};\quad\mathbf{e}_{t}=\mathbf{x}_{t}\oplus\mathbf{r}_{t}$ (3)
where $\mathbf{x}_{t}$ is the question-concept (QC) embedding, $\mathbf{a}^{q}_{t}\in\mathbb{R}^{2\times 1}$ is the one-hot vector indicating whether the question $q_{t}$ is answered correctly and $\mathbf{W}^{a}\in\mathbb{R}^{d\times 2}$ is learnable linear transformation operation. The illustration of the interaction encoding procedure is shown in Figure 2 (a).
4.1.2 Pre-training
Recently, generative pre-trained models that based on Transformer architecture have achieved promising results in various tasks compared to designing a sophisticated neural network for a specific task [38, 13, 14]. Drawing inspiration from these impressive findings, we opt to directly utilize a stack of transformer decoders. This choice enables us to dynamically capture student knowledge states without additional architecture design efforts:
$\mathbf{h}^{(0)}=\mathbf{E}\oplus\mathbf{P}$
$\mathbf{h}^{(l)}=\mathbf{Tranformer\_block}(\mathbf{h}^{(l-1)})\quad\forall{l}\in[1,L]$ (4)
where $\mathbf{E}=(\mathbf{e}_{1},...,\mathbf{e}_{T})$ is the embedding matrix of $T$ past interaction. $\mathbf{P}$ is the position embedding matrix. $L$ is the number of layers. $\mathbf{h}^{(l)}\in\mathbb{R}^{T\times d}$ is a knowledge state embedding matrix of a student by $T$ past interactions. Please note that to estimate student knowledge states via their historical interactions, we use QC embedding $\mathbf{x}_{t}$ for mapping both queries and keys, and interaction embedding $\mathbf{e}_{t}$ for mapping values in the self-attention mechanism.
Furthermore, inspired by prompt learning techniques that effectively capture diverse and overlapping patterns in multi-task learning scenarios [39, 40, 41, 42], we introduce dedicated dataset embeddings for individual KT datasets. Given that each KT dataset has its unique prediction paradigm stemming from variations in question banks and KCs, this enhancement empowers the pre-trained KT model to effectively capture the specific and shared information across different KT datasets. Specifically, the knowledge state $\mathbf{h}^{(l)}$ is first concatenated with corresponding dataset embedding $\mathbf{d}_{i}$ and QC embedding $\mathbf{x}_{t+1}$, then fed into a two-layer fully connected network with Sigmoid activation function $\sigma(\cdot)$ to predict the performance of a student on next question $q_{t+1}$:
$\mathbf{d}_{i}=\mathbf{W}^{d}\cdot\mathbf{e}^{d}$
$\mathbf{y}_{t+1}=\mbox{ReLU}(\mathbf{W}_{1}\cdot[\mathbf{h}_{t+1}^{(l)};\mathbf{x}_{t+1};\mathbf{d}_{i}]+\mathbf{b}_{1})$ (5)
$\hat{r}_{t+1}=\sigma(\mathbf{w}^{\top}\cdot\mbox{ReLU}\bigl{(}\mathbf{W}_{2}\cdot\mathbf{y}_{t+1}+\mathbf{b}_{2}\bigl{)}+b)$
where $\mathbf{W}^{d}\in\mathbb{R}^{d\times I}$ is learnable linear transformation operation [^4] In the fine-tuning stage, we expand the size of $\mathbf{W}^{d}$ to assign the dataset embedding for the specific low-resource KT dataset., $I$ is the total number of rich-resource KT datasets in the pre-training stage. $\mathbf{e}^{d}\in\mathbb{R}^{I\times 1}$ is the one-hot vector indicating the corresponding dataset that current interaction belongs to. $\mathbf{W}_{1}\in\mathbb{R}^{d\times 2d}$, $\mathbf{W}_{2}\in\mathbb{R}^{d\times d}$, $\mathbf{w}\in\mathbb{R}^{d\times 1}$, $\mathbf{b}_{1}\in\mathbb{R}^{d\times 1}$, $\mathbf{b}_{2}\in\mathbb{R}^{d\times 1}$ and $b$ are trainable parameters. All learnable parameters in LoReKT are trained in end-to-end fashion by minimizing the binary cross entropy loss between predicted probability $\hat{r}_{t}$ and the ground-truth label $r_{t}$:
$\mathcal{L}_{\text{KT}}=-\sum_{t=1}^{T}\bigl{(}r_{t}\log\hat{r}_{t}+(1-r_{t})\log(1-\hat{r}_{t})\bigl{)}$ (6)
The forward procedure is illustrated in Figure 2 (b).
Figure 3: An illustration of the importance vector computing and applying procedure in our LoReKT framework.
4.2 Fine-tuning Stage
In low-resource scenarios, overfitting is a common problem, as some model parameters may learn and memorize noisy dataset information, thereby hindering the model’s ability to generalize. This problem is especially severe in the low-resource KT setting. To alleviate the above problem, we introduce a novel importance vector-based fine-tuning strategy to encourage model to focus on updating the important parameters while constraining less important ones.
4.2.1 Computing Importance Vector of Layer
The backbone of LoReKT is a stack of transformer decoders. The key components of a transformer decoder are multi-head attention layer, intermediate layer, and output layer [^5] In this paper, we use “layer” or $l$ to indicate any of these three layers, because the procedure of computing these three layers’ importance vector is similar.. It has been found that not all units (neurons or attention heads) in a specific layer are important [43]. Therefore, before directly fine-tuning the model on each low-resource KT dataset, we adopt the approach described by Ke et al. [44] to compute the importance vector for each layer. This is achieved by employing a gradient-based importance detection method, which is specifically tailored for each low-resource KT dataset:
$\hat{\mathbf{o}}_{l}=\mathbf{g}_{l}\odot\mathbf{o}_{l}$
$\mathbf{I}_{l}=\frac{1}{N}\sum_{n=1}^{N}\left|\frac{\partial\mathcal{L}_{\text{KT}}}{\partial\mathbf{g}_{l}}\right|$ (7)
[…]
4.2.2 Fine-tune with Importance Vector
After obtaining the importance vector $\mathbf{I}_{l}$ for each layer in each low-resource KT dataset using the pre-trained model, we initially compute the original gradient $\nabla_{l}$ by employing $\mathcal{L}_{\text{KT}}$ defined in Eq. 6. Subsequently, we apply the importance vector $\mathbf{I}_{l}$ to obtain the modified gradient ^$\hat{\nabla}_{l}$ for updating:
^$\hat{\nabla}_{l}=\mathbf{I}_{l}\odot\nabla_{l}$
Here, we expand (by copying) the $\mathbf{I}_{l}$ to match the dimensions of $\nabla_{l}$ to apply it to all associated parameters. The modified gradient ^$\hat{\nabla}_{l}$ is only employed in the backward pass. This encourages the model to prioritize updating the associated parameters with high importance instead of less important ones by regulating their gradient flow. The procedure of fine-tuning based on importance vector is shown in Figure 3 (d).
5 Experiment
Low-resource Rich-resource
AS2009 NIPS34 AL2005 BD2006 XES3G5M EdNet
# of Ques. 26,688 948 210,710 207,856 7,652 12,235
# of KCs 123 57 112 493 865 188
# of Interactions 346,860 1,382,727 809,694 3,679,199 5,549,635 6,533,522
avg KCs 1.1969 1.0148 1.3634 1.0136 1.1640 2.2611
Subject Math Math Math Math Math Linguistics
Language English English English English Chinese English
Table 1: Dataset statistics of 6 datasets. “avg KCs” denotes the number of average KCs per question.
In this section, we present details of our experiment settings and the corresponding results. We conduct comprehensive analysis to illustrate the effectiveness of our LoReKT framework. Specifically, we aim to answer the following research questions: (RQ1) Can we build a solid pre-trained foundational model for KT? (RQ2) In low-resource scenarios, how does our proposed LoReKT framework performs compared to the state-of-the-art KT methods? (RQ3) Does pre-training truly enhance the performance of mode in low-resource KT datasets? (RQ4) In the fine-tuning stage, is it effective to focus on updating important parameters based on $\mathbf{I}_{l}$? (RQ5) How does the dataset and data type embedding affect the pre-trained KT model?
5.1 Datasets
Since our LoReKT framework learns transferable parameters and representations from rich-resource KT datasets first and then quickly adapt to low-resource scenarios, we select three rich-resource datasets including BD2006 [47], XES3G5M [48], and EdNet [49] to establish a robust foundational pre-trained KT model. We further fine-tune the pre-trained KT model on three low-resource KT datasets including AS2009 [50], NIPS34 [51], and AL2005 [47] respectively. The data statistics for the six selected datasets can be found in Table 1. The detailed descriptions are as follows:
Rich-resource KT Datasets
  • 1.
    Bridge2algebra2006 (BD2006) [^6] https://pslcdatashop.web.cmu.edu/KDDCup/: this dataset is provided from the KDD Cup 2010 EDM Challenge with the algebra questions answered by 13-14 years old students.
  • 2.
    XES3G5M [^7] https://github.com/ai4ed/XES3G5M: this large-scale dataset from a Chinese online mathematics learning platform includes rich information about student learning interactions.
  • 3.
    EdNet [^8] https://github.com/riiid/ednet: this dataset from South Korea’s Santa AI tutoring system is one of the largest KT datasets, with 130+ million student interactions in the TOEIC test.
Low-resource KT Datasets
  • 1.
    ASSISTments2009 (AS2009) [^9] https://sites.google.com/site/assistmentsdata/home/2009-2010-assistment-data/skill-builder-data-2009-2010: this dataset is one of the classical education dataset that collects students’ responses to mathematic questions from the free online tutoring ASSISTments platform during the school year 2009-2010.
  • 2.
    NIPS34101010https://eedi.com/projects/neurips-education-challenge: this dataset is released in the NeurIPS 2020 Education Challenge. In our work, we choose to use Task 3&4 that are the students’ responses to multiple-choice diagnostic math questions on the Eedi platform.
  • 3.
    Algebra2005 (AL2005) [^11] https://pslcdatashop.web.cmu.edu/KDDCup/: similar to BD2006, this dataset is also released on KDD Cup 2010 EDM Challenge.
5.2 Experimental Setting
We remove student sequences shorter than 3 attempts and truncate student interaction sequences that are longer than 200. We use 80% of student sequences for training and validation and the rest 20% of student sequences for model evaluation. We adopt Adam optimizer [52] to train all the models. The number of training epochs is set to 200. Following all existing DLKT research [15, 24, 36], we use the Area Under the Curve (AUC) as the main evaluation metric and use Accuracy as the secondary evaluation metric. We use an early stopping strategy that stops optimization when the AUC score fails to get the improvement on the validation set in the latest 10 epochs. Owing to the training efficiency, it is difficult to tune too many hyperparameters for the models with billions of parameters. Hence, we only tune the model learning rate in {0.001, 0.0001} with the dropout rate in {0.1, 0.2} for fair comparison in various model sizes.
With the aim of finding a suitable model size for pre-training, we conduct an extensive exploration of various model sizes, encompassing a range from 89M to 1.01B. The corresponding architecture details are summarized in Table 2. $n_{params}$ is the total number of trainable parameters. $n_{layers}$ is the total number of layers, $d_{model}$ is the number of units in each bottleneck layer (the feed-forward layer is denoted as $d_{ff}$), and $n_{head}$ is the number of attention heads.
Model Name $n_{params}$ $n_{layers}$ $d_{model}$ $n_{head}$ $d_{ff}$
LoReKT-Base-89M 89M 4 256 8 256
LoReKT-Base-221M 221M 24 512 16 1024
LoReKT-Base-478M 478M 24 1024 16 1024
LoReKT-Base-1.01B 1.01B 32 1536 24 2560
Table 2: The model sizes and associated architecture details of LoReKT-Base.
5.3 Baselines
To conduct a comprehensive evaluation of LoReKT, we have compared it with 17 selected baseline methods. We have carefully categorized these baseline methods into four categories:
Deep Sequential KT Models
Deep sequential KT models use an auto-regressive framework to dynamically track students’ knowledge states. Representative deep sequential KT models include:
  • 1.
    DKT [16]: it uses an LSTM layer to model students’ learning processes.
  • 2.
    DKT+ [17]: it improves the original DKT model by addressing the reconstruction and inconsistent issues.
  • 3.
    DKT-F [8]: it enhances original DKT by considering students’ forgetting behaviors.
  • 4.
    KQN [18]: it is a recurrent neural network (RNN) based architecture that extracts the relation representations between students’ learning abilities and KCs to predict their performance.
  • 5.
    LPKT [53]: it designs a learning cell to model the students’ learning processes to estimate their knowledge states.
  • 6.
    AT-DKT [54]: it proposes two auxiliary learning tasks involving question tagging prediction task and individualized prior knowledge prediction task to improve the prediction performance of DKT.
Memory Augmented KT Models
Memory augmented KT models employ memory networks to capture potential relevances between KCs and student knowledge states. Representative memory augmented KT models include:
  • 1.
    DKVMN [7]: it incorporates a static matrix to store the relationships among KCs and a dynamic matrix to track the student’s knowledge state.
  • 2.
    SKVMN [55]: it is a combination of DKVMN and LSTM that uses a hop-LSTM layer to capture sequential dependencies of questions.
  • 3.
    DeepIRT [56]: it incorporates DKVMN and item response theory to enhance the interpretability of the prediction output of DKVMN.
Attention based KT Models
Attention based KT models capture dependencies between historical interactions and the next questions via the attention mechanism. Representative attention based KT models include:
  • 1.
    SAKT [19]: it uses self-attention to identify the relevance between historical interactions and KCs.
  • 2.
    SAINT [36]: it is a Transformer-based model for KT that encodes questions and responses in the encoder and decoder respectively.
  • 3.
    AKT [20]: it leverages three self-attention modules to estimate the relevance between questions and historical interactions and explicitly models student’s forgetting behavior via a monotonic attention mechanism.
  • 4.
    simpleKT [24]: it explores the ordinary dot-product attention based KT models by capturing the individual differences among questions covering the same set of KCs
  • 5.
    sparseKT [57]: it incorporates a k-selection module to only pick items with the highest attention scores to improve the robustness and generalization of the attention based DLKT approaches.
Other KT Models
Other KT models that do not belong to the above categories:
  • 1.
    HawkesKT [58]: it utilizes the Hawkes process to model temporal cross-effects in student historical interactions.
  • 2.
    ATKT [10]: it performs adversarial perturbations into the student interaction sequence to enhance the generalization ability based on an attention-LSTM based KT model.
  • 3.
    GKT [26]: it casts the knowledge structure as a graph and reformulates the KT task as a time series node-level classification problem in GNN.
Method (Chronologically) Model Type AUC
AS2009 NIPS34 AL2005 BD2006 XES3G5M EdNet
DKT [16] Sequential 0.7525 0.7688 0.8159 0.8018 0.7845 0.6405
DKVMN [7] Memory 0.7472 0.7677 0.8052 0.7999 0.7796 0.6576
DKT+ [17] Sequential 0.7543 0.7698 0.8141 0.8019 0.7858 0.6454
DKT-F [8] Sequential - 0.7728 0.8146 0.7997 0.7935 0.6548
KQN [18] Sequential 0.7462 0.7685 0.8010 0.7953 0.7794 0.6415
SKVMN [55] Memory 0.7332 0.7513 0.7463 0.7287 0.7512 0.6374
DeepIRT [56] Memory 0.7465 0.7673 0.8040 0.7976 0.7789 0.6387
GKT [26] Others 0.7442 0.7718 0.8112 0.8041 0.7731 0.6392
SAKT [19] Attention 0.7221 0.7508 0.7850 0.7748 0.7685 0.6290
SAINT [36] Attention 0.6990 0.7883 0.7764 0.7758 0.8070 0.6841
AKT [20] Attention 0.7869 0.8038 0.8324 0.8213 0.8215 0.7054
ATKT [10] Others 0.7472 0.7664 0.7987 0.7889 0.7791 0.6490
HawkesKT [58] Others 0.7232 0.7763 0.8199 0.8077 0.7933 0.7304
LPKT [53] Sequential 0.7812 0.8004 0.8268 0.8056 0.8163 0.7644
AT-DKT [54] Sequential 0.7555 0.7816 0.8246 0.8104 0.7925 0.6536
simpleKT [24] Attention 0.7744 0.8035 0.8254 0.8160 0.8161 0.6765
sparseKT [57] Attention 0.7739 0.8033 0.8152 0.8120 0.8165 0.6804
LoReKT-Base-89M Attention 0.6041 0.6401 0.5943 0.8049 0.8145 0.7647
LoReKT-Base-221M Attention 0.6228 0.6452 0.6155 0.8183 0.8192 0.7672
LoReKT-Base-478M Attention 0.5957 0.6103 0.5834 0.8061 0.8164 0.7659
LoReKT-Base-1.01B Attention 0.5761 0.5980 0.5745 0.8003 0.8121 0.7633
LoReKT-Ft-impt-221M Attention 0.7912 0.8002 0.8425 - - -
LoReKT-Ft-221M Attention 0.7833 0.7969 0.8359 - - -
Table 3: The overall performance in terms of AUC. The result of each low-resource KT dataset corresponds to a separately fine-tuned model, leading to different performance on pre-training datasets. Therefore, we use “-” to denote the results on pre-training datasets. The best result is indicated in bold, while the second best result is denoted in underline.
Method (Chronologically) Model Type Accuracy
AS2009 NIPS34 AL2005 BD2006 XES3G5M EdNet
DKT [16] Sequential 0.7228 0.7031 0.8105 0.8554 0.8173 0.6665
DKVMN [7] Memory 0.7196 0.7022 0.8026 0.8547 0.8155 0.6392
DKT+ [17] Sequential 0.7243 0.7046 0.8085 0.8554 0.8179 0.6668
DKT-F [8] Sequential - 0.7076 0.8092 0.8540 0.8209 0.6666
KQN [18] Sequential 0.7214 0.7028 0.8022 0.8540 0.8154 0.6665
SKVMN [55] Memory 0.7156 0.6885 0.7837 0.8406 0.8071 0.6572
DeepIRT [56] Memory 0.7196 0.7020 0.8029 0.8542 0.8152 0.6559
GKT [26] Others 0.7179 0.7053 0.8078 0.8555 0.8139 0.6672
SAKT [19] Attention 0.7031 0.6873 0.7948 0.8460 0.8121 0.6519
SAINT [36] Attention 0.6977 0.7187 0.7770 0.8455 0.8177 0.6624
AKT [20] Attention 0.7385 0.7320 0.8138 0.8594 0.8275 0.6888
ATKT [10] Others 0.7206 0.7012 0.7989 0.8510 0.8143 0.6642
HawkesKT [58] Others 0.7046 0.7110 0.8108 0.8563 0.8191 0.7076
LPKT [53] Sequential 0.7355 0.7309 0.8154 0.8547 0.8264 0.7243
AT-DKT [54] Sequential 0.7250 0.7146 0.8144 0.8560 0.8195 0.6684
simpleKT [24] Attention 0.7320 0.7328 0.8083 0.8579 0.8240 0.6624
sparseKT [57] Attention 0.7282 0.7322 0.8017 0.8569 0.8234 0.6643
LoReKT-Base-89M Attention 0.6035 0.6003 0.6753 0.8543 0.8248 0.7243
LoReKT-Base-221M Attention 0.6216 0.6008 0.6962 0.8596 0.8271 0.7250
LoReKT-Base-478M Attention 0.5929 0.5821 0.6644 0.8538 0.8253 0.7212
LoReKT-Base-1.01B Attention 0.5805 0.5745 0.6568 0.8501 0.8239 0.7207
LoReKT-Ft-impt-221M Attention 0.7402 0.7323 0.8242 - - -
LoReKT-Ft-221M Attention 0.7353 0.7275 0.8159 - - -
Table 4: The overall performance in terms of Accuracy. The result of each low-resource KT dataset corresponds to a separately fine-tuned model, leading to different performance on pre-training datasets. Therefore, we use “-” to denote the results on pre-training datasets. The best result is indicated in bold, while the second best result is denoted in underline.
5.4 Results
We utilize different variants of LoReKT to represent its performance under different settings. LoReKT-Base represents the model trained after the pre-training stage without any fine-tuning on a specific low-resource dataset. LoReKT-Ft-impt and LoReKT-Ft refer to the model that fine-tuned on the specific low-resource KT dataset with and without using importance vector.
5.4.1 Model Performance after Pre-training (RQ1)
We report the results of main evaluation metric, i.e., AUC, in Table 3 and the results of secondary evaluation metric, i.e., Accuracy, in Table 4. From Table 3, we have the following observations: (1) in the comparison of various model sizes in LoReKT-Base, LoReKT-Base-221M exhibits the best performance. Initially, as the model size increases, the model’s performance improves; however, it begins to decline when the model size becomes excessively large. We argue that this phenomenon is indicative of the model first experiencing underfitting followed by overfitting; (2) the LoReKT-Base-221M demonstrates strong performance across all three pre-training datasets (BD2006, XES3G5M, and EdNet). For example, it achieves the highest AUC score on the EdNet dataset, showcasing a substantial improvement of 12.59% over sparseKT, 9.2% over LPKT and 8.6% over AKT. It ranks second in terms of AUC on BD2006 and XES3G5M datasets and is on par with the best model AKT within a 0.5% range of performance gap. It’s noteworthy that LoReKT-Base-221M, utilizing only a single model, consistently achieves strong performance across all three datasets. In contrast, the AKT model is separately trained for each dataset, resulting in significant performance variations. For instance, while the AKT model performs well on BD2006 and XES3G5M, its performance on EdNet is notably lower, lagging behind LoReKT-Base-221M by 8.1%. Furthermore, the architecture of LoReKT-Base-221M is more concise than AKT which designs two extra modules upon the original RNN architecture; and (3) the LoReKT-Base-221M also demonstrates impressive zero-shot capabilities on previously unseen datasets (AS2009, NIPS34, and AL2005). For example, it achieves AUC scores of 0.6452 and 0.6228 in NIPS34 and AS2009 datasets, which is significantly outperforming random performance. In spite of the zero-shot performance is still far from usable, it verifies our conjecture that LoReKT-Base has good potential transferability between different disciplines via cross-source learning. Furthermore, its robust zero-shot capabilities enhance the accuracy of the importance vector computed in Section 4.2.1 for each low-resource KT dataset in the fine-tuning stage.
5.4.2 Fine-tuning Performance of LoReKT (RQ2)
After obtaining the pre-trained KT model LoReKT-Base-221M, we further fine-tune it on each low-resource KT dataset. From Table 3, we can observe that (1) comparing LoReKT-Base-221M and LoReKT-Ft-impt-221M, the performance on all three low-resource KT datasets is significantly improved in terms of AUC score (e.g., an improvement of 26.1% in AS2009, 24.3% in NIPS34 and 37.3% in AL2005); (2) the LoReKT-Ft-impt-221M outperforms all the baseline methods in AS2009 and AL2005 datasets in terms of AUC score. More specifically, compared with ATKT, sparseKT, and simpleKT, it significantly improves the AUC score by 5.61%, 3.47%, and 2.16% on AL2005 respectively. Also, it outperforms the majority of baseline methods and closely matches the performance of the top models such as AKT and simpleKT on NIPS34, with performance gap of less than 0.5%. Notably, our LoReKT-Ft-impt-221M is robust enough to consistently achieve strong performance across all three low-resource datasets without the need for additional architecture design, while AKT and simpleKT exhibit notable performance variations across these datasets. Furthermore, it can be observed that the performance of LoReKT-Ft-impt-221M on NIPS34 is slightly lower than on AS2009 and AL2005 datasets when compared to the baseline methods. We attribute this discrepancy to the larger dataset size of NIPS34 relative to AS2009 and AL2005 (shown in Table 1), which suggests that the potential overfitting issue is not as prominent in NIPS34.
5.4.3 Impact of Pre-training (RQ3)
To analyze the impact of pre-training on the performance of low-resource KT datasets, we progressively reduce the number of rich-resource datasets used in the pre-training. From Figure 4, we have the following observations: (1) As the number of rich-resource KT datasets used in the pre-training stage decreases, the model’s performance on low-resource KT datasets correspondingly drops. (2) When pre-training is omitted, the model’s performance on low-resource KT datasets significantly deteriorates, particularly for the AS2009 and AL2005 datasets. For example, there is a decline in the AUC score by 3.6% in AS2009 and 2.1% in AL2005, whereas only a 1.1% decline is observed in NIPS34. This could be attributed to the fact that the data quantity in AS2009 and AL2005 is much smaller than in NIPS34, leading to a more pronounced overfitting issue.
Figure 4: The impact of the number of rich-resource KT datasets used in the pre-training stage on the performance of low-resource KT datasets.
5.4.4 Impact of Importance Vector (RQ4)
We conduct experiments to further investigate the effectiveness of our proposed fine-tuning strategy with importance vector in low-resource KT datasets. Comparing the results of LoReKT-Ft-221M and LoReKT-Ft-impt-221M in Table 3, we observe that the proposed fine-tuning strategy with importance vector enhances performance across all low-resource datasets. It leads to AUC improvements of 0.66% for the AL2005 dataset and 0.79% for the AS2009 dataset, which is larger than the 0.33% improvement observed for the NIPS34 dataset. This observation suggests that the fine-tuning strategy with importance vector is more effective when the dataset size is insufficient (as shown in Table 1, NIPS34 exhibits a higher interaction count compared to AS2009 and AL2005). We believe that this effectiveness is attributed to the importance vector, which restricts the update of unimportant parameters, thereby mitigating the risk of overfitting.
Method AUC Accuracy
BD2006 XES3G5M EdNet BD2006 XES3G5M EdNet
LoReKT-Base-221M 0.8183 0.8192 0.7672 0.8596 0.8271 0.7250
w/o data type & dataset embedding 0.8145 0.8139 0.7594 0.8541 0.8239 0.7185
w/o data type embedding 0.8161 0.8172 0.7646 0.8575 0.8268 0.7243
w/o dataset embedding 0.8173 0.8168 0.7625 0.8569 0.8268 0.7246
Table 5: The impact of data type and dataset embedding in LoReKT-Base in terms of AUC and Accuracy performance. ”w/o“ means excluding such module from LoReKT-Base. The best result is indicated in bold, while the second best result is denoted in underline.
5.4.5 Impact of Data Type and Dataset Embeddings (RQ5)
We also analyze the impact of data type and dataset embedding in Section 4.1. We report the AUC and Accuracy results in Table 5. As presented in Table 5, comparing the results of “w/o data type & dataset embedding” and “w/o data type embedding”, we demonstrate that the proposed dataset embedding enhances the model’s understanding of the corresponding dataset’s prediction paradigm. This effect is particularly pronounced for datasets with larger quantities, such as the improvement in AUC of 0.52% for EdNet, surpassing the improvements of 0.33% in XES3G5M and 0.16% in BD2006. We attribute this to the fact that the dataset embedding is well-represented when the corresponding dataset has a sufficient quantity (as indicated in Table 1). The results of “w/o data type & dataset embedding” and “w/o dataset embedding” reveal that, across all three datasets, the proposed data type embedding plays an important role in enabling the model to incorporate information from both questions and concepts (e.g., an improvement in AUC of 0.28%, 0.29% and 0.31% in BD2006, XES3G5M, and EdNet respectively).
6 Conclusion
In this paper, we focus on improving the performance of DLKT models on low-resource KT datasets. To address this problem, we propose a framework called LoReKT based on a stack of transformer decoders. The LoReKT includes two stages: pre-training and fine-tuning, and does not require sophisticated architecture for a specific dataset. In the pre-training stage, we establish a robust pre-trained KT model based on several rich-resource KT datasets. Subsequently, we leverage an importance mechanism fine-tuning strategy to adapt the pre-trained model to a specific low-resource KT dataset effectively. The extensive quantitative and qualitative experiment results on six real-world datasets demonstrate the superior performance of LoReKT against a wide range of recently proposed DLKT models in terms of AUC and Accuracy.
CRediT authorship contribution statement
Hengyuan Zhang: Methodology, Conceptualization, Investigation, Writing.
Zitao Liu: Writing - review & editing, Supervision.
Shuyan Huang: Writing - review & editing, Supervision.
Chenming Shang: Investigation, Methodology, Writing.
Bojun Zhan: Investigation, Writing.
Yong Jiang: Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.